# SITE: towards Spatial Intelligence Thorough Evaluation

### v1

@@ Abstract
Spatial intelligence (SI) represents a cognitive ability encompassing the visualization, manipulation, and reasoning about spatial relationships, underpinning disciplines from neuroscience to robotics. We introduce SITE, a benchmark dataset towards SI Thorough Evaluation in a standardized format of multi-choice visual question-answering, designed to assess large vision-language models’ spatial intelligence across diverse visual modalities (single-image, multi-image, and video) and SI factors (figural to environmental scales, spatial visualization and orientation, intrinsic and extrinsic, static and dynamic). Our approach to curating the benchmark combines a bottom-up survey about 31 existing datasets and a top-down strategy drawing upon three classification systems in cognitive science, which prompt us to design two novel types of tasks about view-taking and dynamic scenes. Extensive experiments reveal that leading models fall behind human experts especially in spatial orientation, a fundamental SI factor. Moreover, we demonstrate a positive correlation between a model’s spatial reasoning proficiency and its performance on an embodied AI task. Code and data will be publicly available.
@@ 1 Introduction
This work introduces SITE, a novel benchmark dataset towards Spatial Intelligence Thorough Evaluation, to assess the visuospatial ability of large vision-language models (VLMs). We achieve this by borrowing three classification systems about spatial intelligence (SI) from the cognitive science literature [53, 51, 45, 7] to analyze 8,068 vision-language tasks derived from 31 computer vision datasets. This process highlights a gap in existing benchmarks, leading us to design new tasks focusing on spatial orientation (view-taking) in static and dynamic contexts. We standardize all tasks to ease evaluation using a multiple-choice visual question answering (VQA) format.
@@
SI represents a cognitive capacity encompassing the visualization, manipulation, and reasoning about spatial relationships [20]. Figure SITE: towards Spatial Intelligence Thorough Evaluation shows some SI tests at different scales (left panel), that require spatial visualization independent of one’s viewpoint and spatial orientation due to changes of viewpoints (middle), extrinsic, and dynamic (right). SI is essential for many professions, including architecture, engineering, and the arts, and SI is visual — some works in cognitive science use SI and “visuospatial abilities” interchangeably [60]. While the computer vision community has addressed some components of SI through tasks such as object detection [15, 41, 21, 77, 71], object referring [32, 78, 49, 75], 2.5D visual relationships [61], and counting/localization in VQA [2, 28, 50], progress has occurred mainly within individual domains or datasets. However, the advent of general-purpose large VLMs necessitates a unified testbed incorporating these diverse tasks and some uncaptured aspects of SI. This work presents such a benchmark and a systematic approach to building it.
@@
The role of SI in AI models is highly tied to its role in human perception and reasoning. Kant argued that “space is the form of outer intuition”, asserting that our perception of the physical world is structured by spatial relations between objects [31]. This concept is equally crucial for AI models, which have to develop SI to navigate and interact effectively in complex environments, especially with regard to tasks including but not limited to object manipulation [33, 63, 12, 5] and navigation [23, 3, 59, 27]. In this work, we specifically evaluate large VLMs. These models have demonstrated impressive capabilities in visual reasoning and question answering, thus positioning them as key components for perception and reasoning in embodied AI agents and robotics [33, 6]. However, their ability to perform fine-grained spatial reasoning remains relatively limited and only partially assessed. We expect this work will facilitate a comprehensive and holistic evaluation of VLMs across as broad a spectrum of SI as possible.
@@
SITE (ours) VSI-Bench [74] 3DSRBench [47] CVBench [64] SpatialEval [69]
@@
The proposed SITE benchmark poses a stark contrast to several similar efforts about SI. Table 1 summarizes the key differences. The visual part in SITE combines natural images, synthetic images, multiple views, and videos, while the existing ones contain only natural images [64, 47] or videos [74]. More importantly, these benchmarks measure limited aspects of spatial intelligence. For instance, CVBench [64] lacks viewpoint transformation. 3DSRBench[47] is limited to single-image questions, overlooking spatial reasoning in dynamic contexts. While VSI-Bench [74] uses videos to evaluate VLMs’ capability to perform spatial reasoning across time, it is constrained to only indoor scenes. Finally, while counting and localization frequently appear in VQA benchmarks [2, 28, 44], tasks that require reasoning across multiple viewpoints remain largely unaddressed. These gaps manifest significant challenges in comprehensively evaluating VLMs’ spatial intelligence.
To address the gaps, we approach the curation of SITE from two complementary paths: Bottom-Up and Top-Down. In the bottom-up path, we survey 31 representative datasets and systematically extract 8,068 vision-language tasks after careful filtering. The filtering comprises two phases. We first prompt large language models (LLMs) using the language part of the tasks to reduce costs, and we then filter the surviving tasks by jointly screening their vision-language modalities. Finally, we identify six core categories from the tasks. This bottom-up approach gives rise to 8,068 tasks, including 4,260 image-based QA pairs and 3,808 video-based QA pairs. The top-down approach draws upon three classification systems of SI from the cognitive science literature, capturing SI’s primary factors from different perspectives: scales (figural, vista, and environmental), view-taking (spatial visualization and orientation), intrinsic vs. extrinsic structures, and static vs. dynamic scenes. Investigation shows that the tasks resulting from the bottom-up path underrepresent the view-taking and dynamic factors, so we design two novel types of tasks using the Ego-Exo4D dataset [22], which is rich in camera views of dynamic events. SITE unions the bottom-up and top-down tasks and standardizes them for the ease of evaluations.
@@
    Evident Correlation Between SI and Embodied AI. Finally, we empirically demonstrate that VLMs with high visuospatial ability also perform well on robot manipulation, with a correlation coefficient of 0.922.
@@ 2 Related Work
SI in cognitive science. SI has historically been studied as the interaction of multiple sensory modalities, including vision, touch, and hearing [13]. Out of these modalities, vision has been recognized as the dominant modality facilitating sensory integration [60], making visual-spatial ability the primary focus of mainstream spatial intelligence assessments. Psychologists and cognitive scientists have attempted to define and decompose spatial ability through factor-analytic studies [53, 51, 45, 7, 66, 24] based on numerous paper-and-pencil tests [14]. Several core factors have consistently emerged in these studies, including Spatial Visualization, Spatial Relations, and Spatial Orientation [53, 51, 45]. Later, Carroll et al. [7] introduced an additional factor: Visual Memory. Uttal [66] and Hegarty [24] each proposed new classification models for spatial intelligence. Due to variations in experimental paradigms and analytical methodologies, the definition of spatial ability remains highly inconsistent—a widely acknowledged consensus in cognitive science and psychology [67]. In Section 3, we will further elaborate on these concepts.
@@
SI in computer vision. Spatial visual reasoning has long been an active research topic in computer vision. Initial efforts have mostly focused on constructing large-scale image-based datasets [2, 28, 43, 30], that include spatial visual reasoning tasks, to evaluateVQA approaches. Notable examples such as VQA [2], GQA [28], VSR [43], and CLEVR [30] incorporate spatial reasoning tasks, often evaluating a model’s ability to reason about spatial relationships between objects within an image [2]. However, these datasets generally focus on relatively simple and straightforward tasks, such as verifying the correctness of spatial relationships between objects in an image [2]. This limitation negatively affects their ability to evaluate more complex aspects of spatial intelligence. Meanwhile, advancements in 3D vision and autonomous driving have significantly enriched spatial visual task datasets. Pioneering datasets such as ScanNet [11] and NuScenes [57] provide high-quality 3D annotations, thus facilitating the construction of spatial reasoning benchmarks [8, 4, 57], which leverage multi-image and multi-view inputs to increase spatial task complexity. However, these benchmarks are still largely about static scenes. In contrast, our work on SITE aims to mitigate this limitation by providing a systematic evaluation of spatial intelligence at multiple scales and of both dynamic and static scenes.
@@
Spatial intelligence for benchmarking VLM models. As mentioned earlier, many recent VLM benchmarks [17, 44, 40, 18, 80] have acknowledged the importance of spatial intelligence and included relevant evaluation questions. However, these benchmarks are often limited since spatial reasoning is generally treated as one among many tasks, scattered across other evaluation tasks of comprehension, reasoning and perception, such as OCR and Math Reasoning. Existing image-based VLM benchmarks including MME [17], MMBench [44], SpatialEval [69] and CVBench [64] comprise tasks including but not limited to object counting, localization, and question answering about spatial relationships. The Blink [19] dataset is similar in nature to our SITE where it introduces evaluation tasks that involve spatial reasoning from multiple viewpoints and perspectives. Given the inherent difficulty in equipping VLMs with the capability to perform effective spatial reasoning, Cheng et al. [10] propose a new data curation pipeline that leverages 3D scene annotations as well as a module for integrating depth information into VLMs. Similarly, several video understanding benchmarks for VLMs also include tasks related to spatial reasoning. Notable examples of such datasets include MLVU [84], MVBench [40], and VideoMME [18], which evaluate VLMs using videos of varying durations, ranging from seconds to hour-long videos. However, similar to the image counterparts, these video benchmarks do not systematically isolate spatial intelligence as a core focus of their evaluations. Additionally, recent works such as 3DSRBench [47] and SpatialEval [69], heavily emphasize their evaluations on different aspects of spatial reasoning but are primarily limited to single-image evaluations. Similarly, VSI-Bench [74] incorporates video-based spatial tasks but remains constrained to indoor environments. Despite the contributions of these works, the introduced benchmarks generally do not explicitly evaluate VLMs’ spatial reasoning abilities in a structured and comprehensive manner. In contrast, our SITE aims to bridge this gap by comprehensively evaluating spatial intelligence across multiple aspects and diverse visual context.
@@ 3 What Makes Spatial Intelligence (SI)?
We aim to cover various factors of SI comprehensively, and yet the first challenge we encounter is the lack of a consensus on the categorization of SI [67]. To the best of our understanding, at least three major classification systems about SI exist in cognitive science, and we refer to them all when building our benchmark.
@@
Figural, vista, and environmental SI. Hegarty et al. study SI at figural, vista, and environmental scales [26, 54] and note that “processing spatial information at different scales of space involves different brain structures and mechanisms,” inspiring us to include visual inputs at all these scales. Figural space is small relative to the human body and can be apprehended from a single viewpoint. The following cognitive tests are in this space: mentally folding paper, transforming shapes, and solving mazes. Vista space, such as single rooms and town squares, is larger than the body and remains apprehendable from a single viewpoint. Environmental space contains an individual, referring to cities, neighborhoods, etc., which one must navigate to apprehend. Figure SITE: towards Spatial Intelligence Thorough Evaluation(left) presents examples at the three scales.
@@
Spatial visualization & orientation. Hegarty and Waller compiled a summary of SI factors identified from primary cognitive studies [25], including spatial visualization (VZ), spatial orientation (SO), kinesthetic imagery [53], relational reasoning [45], visual memory [7], and others. Of these, VZ and SO are considered fundamental (see Figure SITE: towards Spatial Intelligence Thorough Evaluation(middle) for examples). Spatial visualization refers to the ability to mentally manipulate, rotate, or invert objects independently of one’s own perspective, while spatial orientation involves imagining object appearances from differing observer viewpoints. It is important to note that these factors are derived from cognitive tests conducted prior to 2000, which primarily assess SI at figural and vista scales. It is also worth noting the extensive suite of cognitive tests about SI designed by Eliot and Ian Macfarlane [14].
@@
2x2 classification. Uttal et al. categorize SI using a 2x2 classification system that relies on two distinctions, intrinsic vs. extrinsic and static vs. dynamic. Figure SITE: towards Spatial Intelligence Thorough Evaluation(right) demonstrates this system. Intrinsic information refers to an object’s defining features, parts, and the relationships among the parts. Extrinsic information, conversely, concerns the spatial relationships between objects within a group or their relation to a broader framework. Static tasks are about fixed spatial information (e.g., counting chairs in a dining room), while dynamic tasks involve movement and transformation (e.g., mental rotation of a 3D shape).
@@
In what follows, we employ all three categorization systems to drive our overarching design complementarily; meanwhile, we rely on them to balance and examine the coverage of the resulting benchmark from different aspects. Utilizing three rather than one categorization marks a significant difference between our work and most existing benchmarking efforts [74, 47, 44, 17, 82, 57, 40].

### v2

@@ Abstract
Spatial intelligence (SI) represents a cognitive ability encompassing the visualization, manipulation, and reasoning about spatial relationships, underpinning disciplines from neuroscience to robotics. We introduce SITE, a benchmark dataset towards SI Thorough Evaluation in a standardized format of multi-choice visual question-answering, designed to assess large vision-language models’ SI across diverse visual modalities (single-image, multi-image, and video) and SI factors (figural to environmental scales, spatial visualization and orientation, intrinsic and extrinsic, static and dynamic). Our approach to curating the benchmark combines a bottom-up survey of existing datasets and a top-down strategy drawing upon three classification systems in cognitive science, which prompt us to design two novel types of tasks about view-taking and dynamic scenes. Extensive experiments reveal that leading models fall behind human experts, especially in spatial orientation, a fundamental SI factor. Moreover, we demonstrate a positive correlation between a model’s spatial reasoning proficiency and its performance on an embodied AI task.
@@ 1 Introduction
This work introduces SITE, a novel benchmark dataset towards Spatial Intelligence Thorough Evaluation, to assess the visuospatial ability of large vision-language models (VLMs). We achieve this by borrowing three classification systems about spatial intelligence (SI) from the cognitive science literature [michael1957description, mcgee1979human, lohman1988spatial, carroll1993human, Kim_2021] to analyze vision-language tasks derived from 30 computer vision datasets. This process highlights a gap in existing benchmarks, leading us to design new tasks focusing on spatial orientation (view-taking) in static and dynamic contexts. We standardize all tasks to ease evaluation using a multiple-choice visual question answering (VQA) format.
@@
SI represents a cognitive capacity encompassing the visualization, manipulation, and reasoning about spatial relationships [gardner2011frames]. Figure 1 shows some SI tests at different scales (left panel), that require spatial visualization independent of one’s viewpoint and spatial orientation due to changes of viewpoints (middle), extrinsic, and dynamic (right). SI is essential for many professions, including architecture, engineering, and the arts, and SI is visual — some works in cognitive science use SI and “visuospatial abilities” interchangeably [shah2005cambridge]. While the computer vision community has addressed some components of SI through tasks such as object detection [Everingham10, lin2014microsoft, geiger2015kitti, yu2020bdd100k, xiang2016objectnet3d], object referring [kazemzadeh2014referitgame, yu2016modeling, mao2016generation, yang2019fast], 2.5D visual relationships [su20222], and counting/localization in VQA [agrawal2016vqavisualquestionanswering, hudson2018gqa, marino2019ok], progress has occurred mainly within individual domains or datasets. However, the advent of general-purpose large VLMs necessitates a unified testbed incorporating these diverse tasks and some uncaptured aspects of SI. This work presents such a benchmark and a systematic approach to building it.
@@
The role of SI in AI models is highly tied to its role in human perception and reasoning. Kant argued that “space is the form of outer intuition", asserting that our perception of the physical world is structured by spatial relations between objects [kant2003critique]. This concept is equally crucial for AI models, which have to develop SI to navigate and interact effectively in complex environments, especially with regard to tasks including but not limited to object manipulation [kim24openvla, team2024octo, ebert2021bridge, bharadhwaj2024roboagent] and navigation [Gu_2022, Anderson_2018_CVPR, shah2023lm, huang2023visual]. In this work, we specifically evaluate large VLMs. These models have demonstrated impressive capabilities in visual reasoning and question answering, thus positioning them as key components for perception and reasoning in embodied AI agents and robotics [kim24openvla, zitkovich2023rt]. However, their ability to perform fine-grained spatial reasoning remains relatively limited and only partially assessed. We expect this work will facilitate a comprehensive and holistic evaluation of VLMs across as broad a spectrum of SI as possible.
@@
SITE (ours) VSI-Bench [yang2024thinkingspacemultimodallarge] 3DSRBench [ma20243dsrbenchcomprehensive3dspatial] CVBench [tong2024cambrian1] SpatialEval [wang2025picture]
@@
The proposed SITE benchmark poses a stark contrast to several similar efforts about SI. Table 1 summarizes the key differences. The visual part in SITE combines natural images, synthetic images, multiple views, and videos, while the existing ones contain only natural images [tong2024cambrian1, ma20243dsrbenchcomprehensive3dspatial] or videos [yang2024thinkingspacemultimodallarge]. More importantly, these benchmarks measure limited aspects of spatial intelligence. For instance, CVBench [tong2024cambrian1] lacks viewpoint transformation. 3DSRBench [ma20243dsrbenchcomprehensive3dspatial] is limited to single-image questions, overlooking spatial reasoning in dynamic contexts. While VSI-Bench [yang2024thinkingspacemultimodallarge] uses videos to evaluate VLMs’ capability to perform spatial reasoning across time, it is constrained to only indoor scenes. Finally, while counting and localization frequently appear in VQA benchmarks [agrawal2016vqavisualquestionanswering, hudson2018gqa, liu2024mmbenchmultimodalmodelallaround], tasks that require reasoning across multiple viewpoints remain largely unaddressed. These gaps manifest significant challenges in comprehensively evaluating VLMs’ spatial intelligence.
To address the gaps, we approach the curation of SITE from two complementary paths: Bottom-Up and Top-Down. In the bottom-up path, we survey 30 representative datasets and systematically extract vision-language tasks after careful filtering. The filtering comprises two phases. We first prompt large language models (LLMs) using the language part of the tasks to reduce costs, and we then filter the surviving tasks by jointly screening their vision-language modalities. Finally, we identify six core categories from the tasks. This bottom-up approach gives rise to 6,943 tasks, including 3,135 image-based QA pairs and 3,808 video-based QA pairs. The top-down approach draws upon three classification systems of SI from the cognitive science literature, capturing SI’s primary factors from different perspectives: scales (figural, vista, and environmental), view-taking (spatial visualization and orientation), intrinsic vs. extrinsic structures, and static vs. dynamic scenes. Investigation shows that the tasks resulting from the bottom-up path underrepresent the view-taking and dynamic factors, so we design two novel types of tasks using the Ego-Exo4D dataset [grauman2024egoexo4dunderstandingskilledhuman], which is rich in camera views of dynamic events. SITE unions the bottom-up and top-down tasks and standardizes them for the ease of evaluations, with a total of 8,068 tasks, covering 30 existing benchmark datasets and 1 newly annotated dataset.
@@
    Evident Correlation Between SI and Embodied AI. Finally, we empirically demonstrate that VLMs with high visuospatial ability also perform well on robot manipulation, with a correlation coefficient of 0.902.
@@ 2 Related Work
SI in cognitive science. SI has historically been studied as the interaction of multiple sensory modalities, including vision, touch, and hearing [Eliot2002]. Out of these modalities, vision has been recognized as the dominant modality facilitating sensory integration [shah2005cambridge], making visual-spatial ability the primary focus of mainstream spatial intelligence assessments. Psychologists and cognitive scientists have attempted to define and decompose spatial ability through factor-analytic studies [michael1957description, mcgee1979human, lohman1988spatial, carroll1993human, Uttal_Meadow_Tipton_Hand_Alden_Warren_Newcombe_2012, Hegarty_2004] based on numerous paper-and-pencil tests [eliot1983international]. Several core factors have consistently emerged in these studies, including Spatial Visualization, Spatial Relations, and Spatial Orientation [michael1957description, mcgee1979human, lohman1988spatial]. Later, Carroll et al. [carroll1993human] introduced an additional factor: Visual Memory. Uttal et al. [Uttal_Meadow_Tipton_Hand_Alden_Warren_Newcombe_2012] and Hegarty and Waller [Hegarty_2004] proposed new classification models for spatial intelligence. Due to variations in experimental paradigms and analytical methodologies, the definition of spatial ability remains highly inconsistent—a widely acknowledged consensus in cognitive science and psychology [uttal2013malleability]. In Section 3, we will further elaborate on these concepts.
@@
SI in computer vision. Spatial visual reasoning has long been an active research topic in computer vision. Initial efforts have mostly focused on constructing large-scale image-based datasets [agrawal2016vqavisualquestionanswering, hudson2018gqa, Liu2022VisualSR, johnson2016clevrdiagnosticdatasetcompositional], that include spatial visual reasoning tasks, to evaluateVQA approaches. Notable examples such as VQA [agrawal2016vqavisualquestionanswering], GQA [hudson2018gqa], VSR [Liu2022VisualSR], and CLEVR [johnson2016clevrdiagnosticdatasetcompositional] incorporate spatial reasoning tasks, often evaluating a model’s ability to reason about spatial relationships between objects within an image [agrawal2016vqavisualquestionanswering]. However, these datasets generally focus on relatively simple and straightforward tasks, such as verifying the correctness of spatial relationships between objects in an image [agrawal2016vqavisualquestionanswering]. This limitation negatively affects their ability to evaluate more complex aspects of spatial intelligence. Meanwhile, advancements in 3D vision and autonomous driving have significantly enriched spatial visual task datasets. Pioneering datasets such as ScanNet [dai2017scannet] and NuScenes [qian2024nuscenesqamultimodalvisualquestion] provide high-quality 3D annotations, thus facilitating the construction of spatial reasoning benchmarks [chen2020scanrefer, azuma_2022_CVPR, qian2024nuscenesqamultimodalvisualquestion, omni3dbench], which leverage multi-image and multi-view inputs to increase spatial task complexity. However, these benchmarks are still largely about static scenes. In contrast, our work on SITE aims to mitigate this limitation by providing a systematic evaluation of spatial intelligence at multiple scales and of both dynamic and static scenes.
@@
SI for benchmarking VLM models. As mentioned earlier, many recent VLM benchmarks [fu2024mmecomprehensiveevaluationbenchmark, liu2024mmbenchmultimodalmodelallaround, li2024mvbenchcomprehensivemultimodalvideo, fu2024video, yu2024mm] have acknowledged the importance of spatial intelligence and included relevant evaluation questions. However, these benchmarks are often limited since spatial reasoning is generally treated as one among many tasks, scattered across other evaluation tasks of comprehension, reasoning, and perception, such as OCR and Math Reasoning. Existing image-based VLM benchmarks including MME [fu2024mmecomprehensiveevaluationbenchmark], MMBench [liu2024mmbenchmultimodalmodelallaround], SpatialEval [wang2025picture] and CVBench [tong2024cambrian1] comprise tasks including but not limited to object counting, localization, and question answering about spatial relationships. The Blink [fu2024blink] dataset is similar in nature to our SITE where it introduces evaluation tasks that involve spatial reasoning from multiple viewpoints and perspectives. Given the inherent difficulty in equipping VLMs with the capability to perform effective spatial reasoning, Cheng et al[cheng2024spatialrgpt] propose a new data curation pipeline that leverages 3D scene annotations as well as a module for integrating depth information into VLMs. Similarly, several video understanding benchmarks for VLMs also include tasks related to spatial reasoning. Notable examples of such datasets include MLVU [MLVU], MVBench [li2024mvbenchcomprehensivemultimodalvideo], and VideoMME [fu2024video]. However, similar to the image counterparts, these video benchmarks do not systematically isolate spatial intelligence as a core focus of their evaluations. Additionally, recent works such as 3DSRBench [ma20243dsrbenchcomprehensive3dspatial] and SpatialEval [wang2025picture], heavily emphasize their evaluations on different aspects of spatial reasoning but are primarily limited to single-image evaluations. Similarly, VSI-Bench [yang2024thinkingspacemultimodallarge] incorporates video-based spatial tasks but remains constrained to indoor environments. Despite the contributions of these works, the introduced benchmarks generally do not explicitly evaluate VLMs’ spatial reasoning abilities in a structured and comprehensive manner while our SITE aims to bridge this gap by evaluating spatial intelligence across multiple aspects and diverse visual context.
@@ 3 What Makes Spatial Intelligence (SI)?
We aim to cover various factors of SI comprehensively, and yet the first challenge we encounter is the lack of a consensus on the categorization of SI [uttal2013malleability]. To the best of our understanding, at least three major classification systems about SI exist in cognitive science, and we refer to them all when building our benchmark.
@@
Figural, vista, and environmental SI. Hegarty et al. study SI at figural, vista, and environmental scales [Hegarty_Montello_Richardson_Ishikawa_Lovelace_2006, montello1993scale] and note that “processing spatial information at different scales of space involves different brain structures and mechanisms,” inspiring us to include visual inputs at all these scales. Figural space is small relative to the human body and can be apprehended from a single viewpoint. The following cognitive tests are in this space: mentally folding paper, transforming shapes, and solving mazes. Vista space, such as single rooms and town squares, is larger than the body and remains apprehensible from a single viewpoint. Environmental space contains an individual, referring to cities, neighborhoods, etc., which one must navigate to apprehend. Figure 1(left) presents examples at the three scales.
@@
Spatial visualization & orientation. Hegarty and Waller compiled a summary of SI factors identified from primary cognitive studies [hegarty2005individual], including spatial visualization (VZ), spatial orientation (SO), kinesthetic imagery [michael1957description], relational reasoning [lohman1988spatial], visual memory [carroll1993human], and others. Of these, VZ and SO are considered fundamental (see Figure 1(middle) for examples). Spatial visualization refers to the ability to mentally manipulate, rotate, or invert objects independently of one’s own perspective, while spatial orientation involves imagining object appearances from differing observer viewpoints. It is important to note that these factors are derived from cognitive tests conducted prior to 2000, which primarily assess SI at figural and vista scales. It is also worth noting the extensive suite of cognitive tests about SI designed by Eliot and Smith [eliot1983international].
@@
2x2 classification. Uttal et al. categorize SI using a 2x2 classification system that relies on two distinctions, intrinsic vs. extrinsic and static vs. dynamic. Figure 1(right) demonstrates this system. Intrinsic information refers to an object’s defining features, parts, and the relationships among the parts. Extrinsic information, conversely, concerns the spatial relationships between objects within a group or their relation to a broader framework. Static tasks are about fixed spatial information (e.g., counting chairs in a dining room), while dynamic tasks involve movement and transformation (e.g., mental rotation of a 3D shape).
@@
In what follows, we employ all three categorization systems to drive our overarching design complementarily; meanwhile, we rely on them to balance and examine the coverage of the resulting benchmark from different aspects. Utilizing three rather than one categorization marks a significant difference between our work and most existing benchmarking efforts [yang2024thinkingspacemultimodallarge, ma20243dsrbenchcomprehensive3dspatial, liu2024mmbenchmultimodalmodelallaround, fu2024mmecomprehensiveevaluationbenchmark, lmms_eval2024, qian2024nuscenesqamultimodalvisualquestion, li2024mvbenchcomprehensivemultimodalvideo].